OpenStudy (loser66):

Let \(f(x,y ) =\dfrac{x+y}{32}\), x =1,2; y = 1,2,3,4 find \(P (1\leq Y\leq 3| X =1)\) Please, help

OpenStudy (loser66):

I have P (X =1) = 14/32

OpenStudy (loser66):

need help on the first part, how to interpret it?

OpenStudy (loser66):

Don't we have to apply Baye's theorem here?

OpenStudy (loser66):

I mean:\(P(Y=1|X=1) =\dfrac{P(Y=1,X=1)}{P(X=1)}\)

OpenStudy (tkhunny):

Yes.

OpenStudy (loser66):

If it is so, then, the first one is (2/32 )/(14/32)= 1/7 different from your answer

OpenStudy (loser66):

So, we just add them up?

OpenStudy (tkhunny):

Yup. I get P(Y=1|X=1) = 1/7

OpenStudy (loser66):

Yes, I got it, thank you so much :)

OpenStudy (tkhunny):

Very confusing typos. I'll delete that.

OpenStudy (loser66):

One more question: how can you get it quickly like that? I have to calculate the probability and fill up the table, calculate everything step by step while you got it right after you see it. Is there any tips?

OpenStudy (tkhunny):

Not really. Use all the tools available to you.

OpenStudy (loser66):

OH,,, if it is so, thanks but I can't. On test, we are not allowed to use technology. :)

OpenStudy (tkhunny):

Well, spread the table and fill it in! Did you also check to see if it actually qualified as a Probability Distribution?

OpenStudy (loser66):

What do you mean ?

OpenStudy (tkhunny):

Does \(\sum{f(Everything)} = 1\)?